RC meeting and Run Plan, Sunday Oct. 1, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary

Despite beam recovery being expected during DAY shift yesterday, production running at 35uA did not restart until almost the end of SWING shift. Some issues with DAQ communication and the SBS GEM HVs stopped the flow of production for 2 hours, but things were smoothed out by 3 am. Constant production runs have been taken since that point, but it was only noticed at 9 that the wrong BBCAL HV settings were being used. In the rush to get beam back into the hall, I had forgotten that the HV settings had been left at the relevant point for BB OFF/SBS OFF since SWING shift on the 28th, and that they needed to be changed since the magnets were now both on.

This issue has been rectified, and production runs are now being taken with the correct HV settings. In addition, the beam was judged to have been stable enough for long enough that, at 1:30 pm, we requested that the beam current be increased to 40uA. Since that point, the beam has been choppy, so an eye will have to be kept on it to see if we need to change current again.

Accelerator status

  • Beam recovery was initiated during DAY shift, but there were problems when trying to run on the C-hole until late into SWING.
  • Beam was restored to the Hall at 11pm, at 35uA.
  • I asked the shift crew to call me if the beam was stable for 2 hours so that we could increase to 40uA; this was judged not to have happened until 1:30 pm today. We are now running at 40uA.
  • No maintenance day on the MCC white board (until 10/10/2023).

Data collected :

  • Accumulated corrected charge collected:
    • 0.005C Day (9/30)
    • 0C Swing (9/30)
    • 0.375C Owl (10/1) - albeit for wrong BBCAL HV.
    • 0.498C Day (10/1)
  • Total corrected charge on cell since 9/27 = 0.920C

Hall A Subsystems Status


  • Latest target polarization: 39.9% [1:54 pm]


  • BBHodo: channel 48R has now got connection issues to the crate. That now makes 2 channels next to each other (48 and 49R), so these should be addressed as soon as an access is available.

Hall access requests

  • BBHodo: investigate 48/49R issues at crate, and 32L low occupancy.
  • If anyone wants to make an access, let me know! There is no planned maintenance day before Tuesday 10th, so any accesses need to be opportunistic.

Shift schedule

  • Still need to fill open slots
  • Open slots on Wednesday 4th: OWL and SWING (both TO).


Please sign up for any open shifts!

Short Term Run plan

  • Beam recovery to the hall
  • Production run on polarized He3 target @35uA beam current.
    • If the beam and target are stable at 35uA, call the RC. We can then go up to 40uA.
    • If 40uA is stable for a few hours, we can consider going up to 45uA - call the RC first!

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)